# Thoughtbubbles: an Unsupervised Method for Parallel Thinking in Latent Space

### v1

@@ 5 Analysis
Figure 4: Normalized number of forks in the final layer across a window of 44 tokens as a function of the mean entropy of those $4$ tokens on OpenWebText . Left: entropy as measured by the forking transformer; right: entropy as measured by a baseline decoder LM.
Forks meaningfully influence the value of the parent token. In fig. 5, we see the rightmost (“og”) token attends to its children with attention scores more than an order of magnitude higher than other tokens—second only to attention of those tokens to themselves. This result indicate that the forking tokens play a large role in the computation of the residual update for the rightmost token than most other tokens, indicating their utility in computing the final output.
Our model dynamically allocates more computation at regions of higher uncertainty without explicit supervision …
Despite no explicit interventions or regularization, our method learned to allocate more computation at areas of greater uncertainty. We see in fig. 4 that our method allocates more computation tokens with high output distribution entropy; this is true both for the entropy measured from the forking model as well as an independently trained, parameter matched decoder LM that does not fork. This is in-line with recent literature (Wang et al., 2025) that highlights the informativeness of high entropy tokens.
@@ 5.1 Autoregression
As seen in fig. 6, implementing autoregression naively with a fixed block size irrespective of the input sequence length results in a distribution shift between blockwise forward pass and autoregression—since the maximum allowed number of forks is much higher if input sequence length is smaller while the total budget remains the same.
However, if we apply the forking budget scaling mitigation described in section E.1, we find that our model performs roughly equivalently to the blockwise forward pass, and retains our approach’s performance gains above baseline. This result indicates that, while our result can adapt to different inference-time input sizes, care must be taken to scale the adaptive computation budget accordingly.
@@ 6 Related Work
Chain-of-thought (Wei et al., 2023) is a simple form of adaptive computation which uses natural-language-based autoregression with additional tokens to achieve thinking. Variants of this approach include simply supervising the output chain (Zhang & Ding, 2024), to replacing them with continuous traces (Hao et al., 2024) or controlled non-adaptive filler tokens (Pfau et al., 2024). Unlike chain-of-thought, our method performs adaptive computation not with recurrence but parallel computation, improving efficiency as well as being able to be trained without additional supervision.
@@
Methods vary to force a dynamic amount of computation from a neural model based on the problem. The oldest approaches involves explicitly forcing recurrence (Graves, 2016), while modern LMs yield performance improvements through forcing very simple interventions to existing chain of thoughts (Muennighoff et al., 2025), skipping or adding recurrent compute across layers without adding additional streams of computation (Dehghani et al., 2019; Csordás et al., 2024; Chen et al., 2024; Murty et al., 2023; Raposo et al., 2024; Kallini et al., 2024), or—most similar to our approach—by adding additional residual streams when computation is needed (Herel & Mikolov, 2024; Goyal et al., 2024; Sun et al., 2025). Unlike prior art, our method removes the need to insert latent tokens explicitly during training or inference, but still gives the ability to gain additional streams of computation through latent residual streams.
@@
There’s a large and robust literature on the complexity-theoretic power of transformers. Results have shown the limited expressive power of standard transformer computation (Merrill & Sabharwal, 2023), and the additional power that chain-of-thought or even padding tokens add to the computation (Merrill & Sabharwal, 2025; London & Kanade, 2025). Work has also shown the limitations given by single special-token thinking approaches that are not input adaptive (Vennam et al., 2024). Prior work have also shown through techniques in intepretability that even simple chain of thought computation carries implicit intermediate computation similar to depth-bounded recurrence (Brinkmann et al., 2024). We also demonstrate here the power of adaptive latent computation in our work by demonstrating its superior performance even against computation matched baselines; furthermore, we demonstrate that we are indeed performing additional computation in “decisive” high entropy tokens, in line with prior analyses (Wang et al., 2025).
@@ 7 Conclusion
This method unlocks the previously missing input-adaptivity of transformer computation, which allows our model to solve more difficult tasks that require scaling inference-time computation. We demonstrate the efficacy of our method via a suite of zero-shot evaluations on models pretrained on both OpenWebText and peS2o in both computation and parameter matched settings. Excitingly, our method at a smaller 319M scale outperformed baselines at 772M scale.
Most importantly, our model enables learning latent adaptive computation in a language model already during the pre-training phase. Unlike CoT approaches, it does not rely on being exposed to step-wise instructions during pre-training.
We hope that this will unlock a new generation of transformer architectures with more general latent computation, which in turn enables more helpful and capable reasoning models.
8 Limitations
Time-matched evaluations
Our current approach is implemented in raw PyTorch, requiring no other hardware-level adaptations apart from scatter-max kernels. [^3] https://github.com/rusty1s/pytorch_scatter As such, though our implementation exceeds the performance of a computation-matched approach, its raw wall-clock efficiency is relatively low. Further efforts in implementing hardware-adaptive kernels for operations like forking gather will enable faster computational performance.
Top-K Gradient Bottleneck
While forking in earlier layers improves the performance of our models, we note that, interestingly, too much forking results in no further performance improvement at a fixed block size budget (appendix B). We believe this is due to certain tokens with high cumulative scores early on in the model being dropped by hard top-k decisions later in the model, thus resulting in no gradients to update the early large cumulative scores. By implementing training time randomization and noise, this can be mitigated to improve deep forking performance.
Downstream Reasoning Tasks
Given our hardware limitations, we cannot experiment with approaches at sufficient scale to enable non-noisy measurements on hard reasoning datasets such as GSM8k (Cobbe et al., 2021), which—without customized training regimes—usually emerges with good performance only around the multi-billion-parameter scale (Liu et al., 2023). In future work, we hope to perform these evaluations with additional resources.
Acknowledgments
We thank Mykel Kochenderfer, Christopher Potts, Tatsunori Hashimoto, Percy Liang, Diyi Yang, Jacob Andreas, and David Bau for their insightful comments and discussion of this project. In addition, we thank our our colleagues for their many interesting discussions, especially Julie Kallini, Dilara Soylu, Steven Cao, Amelia Hardy, Liam Kruse, Sydney Katz, Zachary Sayyah, Huxley Marvit, Joseph Shetaye, and Eric Alcaide. Christopher D.  Manning is a CIFAR Fellow.
Use of Large Language Models
We use LLMs for copyediting only. All ideas and content contained herein are our own.
@@ Appendix A Exact Benchmark Model Architecture
Batch size (global) 64
Steps (global) 75000
@@
Optimization of each run is conducted on a single NVIDIA H200 GPU. Dataset tokenization uses the pre-trained tokenizer from GPT-2 (Radford et al., 2019). FlashAttention kernels (Dao et al., 2022) are used to train our system, with value vector attenuation occurring before.
Appendix B Overforking
We perform a minimal ablation to examine if additional layers of forking beyond layers 3, 7, and 13 would help model performance. In particular, we trained our 772M scale model for 25,000 steps with forking at layers 3, 7, and 11 only as well as extended into all layers 4n14n-1 (i.e., 16, 20, …) thereafter.
@@ Appendix B Overforking
Table 4: Performance of an ablation where we performed more forking at later layers, trained on OpenWebText for 25,000 steps (roughly 0.8BT). We see that the extended forking approach is slightly worse than forking only in the beginning.
Appendix C Analysis of Forking Locations
@@
In fig. 7, we qualitatively observe that our model has learned to perform extra computation near intepretable decision boundaries for synthetic tasks. For CLUTTR, forking occurs near coreferent entities and at special tokens delineating the beginning of query components or response. In contrast, the result for OpenWebText shows that computation on web-text is spread evenly across the sequence—namely, that its not sequence position dependent, as it is for synthetic tasks such as CLUTTR due to their structure.
This result, along with the result in fig. 4, indicates that our model can truly dynamically allocate computation to the areas of the biggest greatest computational difficulty and is not relying on a simple heuristic for allocating computation.
Appendix D Position Encoding
@@
where $\theta$ is the total rotation angle. In our approach, however, we may have $q$ streams representing a particular input token. That is, token $k$ is forked into residual streams $x_{q,k},\dots,x_{0,k}$. In order to accommodate tokens of different number of forks, we augment RoPE with partial rotations proportional to the number of forks of each token. For the $i,j$-th slot of residual $p$ of token $k$ which has $q$ forks in total, we write:
@@
Appendix E Details on Zero-Shot Evals
@@
E.1 Inference-Time Budget Scaling

### v2

@@ 5 Analysis
Here we describe a series of analysis of our approach at the 772M scale to understand its capabilities and behaviors. We further describe ablations to each component of our approach in Appendix C, in particular to understand the token merging behavior (Table 6) and our attention masking approach (Table 7). Finally, we measure and describe our wall-clock performance at (Appendix B), noting that our performance gains are also competitive to timing-based measurements.
Forks meaningfully influence the value of the parent token. In Figure 4, we see the rightmost (“og”) token attends to its children with attention scores more than an order of magnitude higher than other tokens—second only to attention of those tokens to themselves. This result indicate that the forking tokens play a large role in the computation of the residual update for the rightmost token than most other tokens, indicating their utility in computing the final output.
@@
Figure 5: Normalized number of forks in the final layer across a window of 44 tokens as a function of the mean entropy of those $4$ tokens on OpenWebText . Left: entropy as measured by the forking transformer; right: entropy as measured by a baseline decoder LM. Note that our model independently induced the entropy ratings of an unrelated decoder LM.
Our model allocates more computation at regions of higher uncertainty without explicit supervision …
Despite no explicit interventions or regularization, our method learned to allocate more computation at areas of greater uncertainty. We see in Figure 5 that our method allocates more tokens with high output entropy; this is true both for the entropy measured from the forking model as well as an independently trained, parameter-matched decoder LM that does not fork. This is in line with recent literature (Wang et al., 2025) that highlights the informativeness of high entropy tokens.
@@ 5.1 Autoregression
As seen in Figure 6, implementing autoregression naively with a fixed block size irrespective of the input sequence length results in a distribution shift between blockwise forward pass and autoregression—since the maximum allowed number of forks is much higher if input sequence length is smaller while the total budget remains the same.
However, if we apply the forking budget scaling mitigation described in Section G.1, we find that our model performs roughly equivalently to the blockwise forward pass, and retains our approach’s performance gains above baseline. This result indicates that, while our result can adapt to different inference-time input sizes, care must be taken to scale the adaptive computation budget accordingly.
@@ 6 Related Work
Chain-of-thought (Wei et al., 2023) is a simple form of adaptive computation which uses natural-language-based autoregression with additional tokens to achieve thinking. Variants of this approach include simply supervising the output chain (Zhang and Ding, 2024), to replacing them with continuous traces (Hao et al., 2024; Zelikman et al., 2024) or controlled non-adaptive filler tokens (Pfau et al., 2024). Unlike chain-of-thought, our method performs adaptive computation not with recurrence but parallel computation, improving efficiency as well as being able to be trained without additional supervision.
@@
Methods vary to force a dynamic amount of computation from a neural model based on the problem. The oldest approaches involves explicitly forcing recurrence (Graves, 2016), while modern LMs yield performance improvements through forcing very simple interventions to existing chain of thoughts (Muennighoff et al., 2025), skipping or adding recurrent compute across layers without adding additional streams of computation (Dehghani et al., 2019; Murty et al., 2023; Csordás et al., 2024; Chen et al., 2024; Raposo et al., 2024; Kallini et al., 2024), or by adding additional residual streams when computation is needed (Herel and Mikolov, 2024; Goyal et al., 2024; Sun et al., 2025). Our method removes the need to insert latent tokens explicitly during training or inference, but still gives the ability to gain additional streams of computation through latent residual streams.
@@
There’s a large and robust literature on the complexity-theoretic power of transformers. Results have shown the limited expressive power of standard transformer computation (Merrill and Sabharwal, 2023), and the additional power that chain-of-thought or even padding tokens add to the computation (Merrill and Sabharwal, 2025; London and Kanade, 2025). Work has also shown the limitations given by single special-token thinking approaches that are not input adaptive (Vennam et al., 2024). Prior work have also shown through techniques in intepretability that even simple chain of thought computation carries implicit intermediate computation similar to depth-bounded recurrence (Brinkmann et al., 2024). We also demonstrate here the power of adaptive latent computation in our work by demonstrating its superior performance even against computation matched baselines; furthermore, we demonstrate that we are indeed performing additional computation in “decisive” high entropy tokens, in line with prior analyses (Wang et al., 2025).
@@ 7 Conclusion
This method unlocks a form of input-adaptivity in transformer computation, which allows our model to solve more difficult tasks that require scaling inference-time computation. We demonstrate the efficacy of our method via a suite of zero-shot evaluations as well as gsm8k evaluations in both computation and parameter matched settings at 1.9B parameters; furthermore, in our scaling experiments between 150M-772M parameters, we discovered that our method at a smaller 319M scale outperformed baselines at 772M scale.
Most importantly, our model enables learning latent adaptive computation in a language model already during the pre-training phase. Unlike CoT approaches, it does not rely on being exposed to step-wise instructions during pre-training. We hope that this will unlock a new generation of transformer architectures with more general latent computation—a step towards enhancing input-adaptive methods like chain-of-thoughts to be fully latent and integrated as a part of pretraining.
Impact Statement
This paper presents work aimed at advancing the state of the art in neural network architectures. There are many potential societal consequences of our work that come with any general improvements to neural network capability, none of which we believe must be specifically highlighted here.
@@ Appendix A Exact Benchmark Model Architecture
Batch size (global) 64 (150M-772M); 480 (1.9B)
Tokens 2.5B (150-772M); 40B (1.9B)
@@ 1.9B Ours ($\kappa=2L$) 36 16 2048 14 32 1024
Optimization of each run at 150M-772M scaleis conducted on a single NVIDIA H200 GPU. Optimization of each run at 1.9B scale is conducted on a single v4-32 TPU pod. Dataset tokenization uses the pre-trained tokenizer from GPT-2 (Radford et al., 2019). FlashAttention kernels (Dao et al., 2022) and XLA attention kernels, where appropriate, are used to train our system, with value vector attenuation occurring before.
Appendix B Training and Inference Wall-Clock Efficiency
We find that, with an efficient implementation of our approach, and in particular the optimization of residual averaging described in Section 2.5, our forking mechanism is extremely lightweight in terms of wall-clock efficiency—almost matching those of a baseline transformer.
In particular, using torch.compile graph lowering and sequentially blocking CUDA operations (to gain the most accurate timing signals), we obtain the following wall clock performance speeds for a single forward-backward pass of $8$ sequences:
Method Time (ms)
Baseline 234
Ours ($\kappa=2L$) 327
Baseline (2L) 380
Table 5: Wall-clock performance on a single Nvidia H100, with blocking CUDA kernel launch operations, for our approach at $1.9B$ scale and on a batch of 88 sequences; $L=512$ is the block size used.
We find that our full forward/backward pass chain almost matches the efficiency of a baseline transformer, and in particular is faster than a model trained with double the block size. These results, combined with the fact that our approach outperforms baseline at only half of the token count (Table 1), implies competitive wall-clock performance advantage of our approach as well.
Appendix C Gradient Signal Ablations
Use of Forked Tokens’ Gradient Signals
We perform a minimal ablation on our 150M model to analyze whether our model is actually apply gradients to all forking tokens (i.e., instead of ignoring them); in Table 6, we perform a minimal ablation where we keep only one, namely the rightmost, residual channel per input token instead of averaging all forked tokens. We find that this performs worse that our approach.
Approach Number of Tokens (OpenWebText) Validation Loss
Baseline 4.8B 3.27
Ours ($\kappa=2L$, logit avg.) 2.4B 3.17
Ours ($\kappa=2L$, keep rightmost) 2.4B 3.21
Table 6: Performance of an ablation where we keep the rightmost token instead of logit average for output token merging.
Use of Attention Masking Gradient Signals
We additionally perform an ablation to analyze whether our model is using the gradient signal from masking the $QK$ as well as $V$ in attention computation.
Approach Number of Tokens (OpenWebText) Validation Loss
Baseline 4.8B 3.27
Ours ($\kappa=2L$, logit avg.) 2.4B 3.17
Ours ($\kappa=2L$, don’t mask attn) 2.4B 3.90
Table 7: Ablation: removing attention masking significantly degrades performance.
We believe this degradation is due to the fact that cumulative scores are multiplicative; without them affecting computation in some way, the multiplicative scores across layers would simply decrease throughout computation. Thus, the top-$k$ decision is essentially random, resulting in random forks and thus significantly decreased performance.
Appendix D Overforking
We perform a minimal ablation to examine if additional layers of forking beyond layers 3, 7, and 13 would help model performance. In particular, we trained our 772M scale model for 25,000 steps with forking at layers 3, 7, and 11 only as well as extended into all layers 4n14n-1 (i.e., 16, 20, …) thereafter. We find that forking more confers only a slight advantage to performance; we believe this is due to certain tokens with high cumulative scores early on in the model being dropped by hard top-k decisions later in the model, thus resulting in no gradients to update the early large cumulative scores. By implementing training time randomization and noise, this can be mitigated to improve deep forking performance.
@@ Appendix D Overforking
Table 8: Performance of an ablation where we performed more forking at later layers, trained on OpenWebText for 25,000 steps (roughly 0.8BT). We see that the extended forking approach is only slightly better than forking only in the beginning.
Appendix E Analysis of Forking Locations
h Refer to caption Refer to caption
@@ Appendix E Analysis of Forking Locations
In Figure 7, we qualitatively observe that our model has learned to perform extra computation near intepretable decision boundaries for synthetic tasks. For CLUTTR, forking occurs near coreferent entities and at special tokens delineating the beginning of query components or response. In contrast, the result for OpenWebText shows that computation on web-text is spread evenly across the sequence—namely, that its not sequence position dependent, as it is for synthetic tasks such as CLUTTR due to their structure.
This result, along with the result in Figure 5, indicates that our model can truly dynamically allocate computation to the areas of the biggest greatest computational difficulty and is not relying on a simple heuristic for allocating computation.
Appendix F Position Encoding
@@
where $\theta$ is the total rotation angle. In our approach, however, we may have $q$ streams representing a particular input token. That is, token $k$ is forked into residual streams $x_{(q-1),k},\dots,x_{0,k}$. In order to accommodate tokens of different number of forks, we augment RoPE with partial rotations proportional to the number of forks of each token. For the $i,j$-th slot of residual $p$ of token $k$ which has $q$ forks in total, we write:
@@
Appendix G Details on Zero-Shot Evals
@@
AI2-ARC
To evaluate our model’s general in-context reasoning abilities, we measure our model’s performance across both the easy and challenge slices of AI2-ARC (Clark et al., 2018). The dataset contains 4 choices of possible continuations of a given premise, only one of which is correct. A task is solved correctly if the model assigns lower perplexity to the correct continuation conditioned on the premise.
@@
G.1 Inference-Time Budget Scaling